# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

### v1

@@ Abstract
With the exponential growth of video traffic, traditional video streaming systems are approaching their limits in compression efficiency and communication capacity. To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive novel system that streaming prompts instead of video content with Stable Diffusion, which converts video frames into a series of "prompts" for delivery. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. To the best of our knowledge, Promptus is the first attempt to replace video codecs with prompt inversion and the first to use prompt streaming instead of video streaming. Our work opens up a new paradigm for efficient video communication beyond the Shannon limit.
@@ 1 Introduction
With the rapid development of streaming applications (such as video-on-demand [2, 8], live video [14, 7], video conferencing [15, 11], cloud gaming [3, 13], etc.), the traffic of network video has been continuously growing. To reduce traffic, video codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5] are widely used to compress videos. These codecs achieve compression by removing spatial and temporal redundancies. However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit [40]). In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts.footnotetext: Promptus will be released at https://pku-netvideo.github.io/ once published.
@@
Figure 1: Promptus can invert a given image into a prompt. Based on this prompt, Stable Diffusion can generate an almost identical image to the original. In contrast, existing methods can only generate semantically similar images.
To improve compression efficiency, some deep learning-based codecs and streaming frameworks have been proposed in recent years. Neural-embedded codecs [31, 30, 19, 37, 47] replace handcrafted modules (such as motion vector estimation [45]) in traditional codec frameworks with neural networks (such as autoencoders), learning more intelligent and efficient compression capabilities through end-to-end optimization. However, due to still following the traditional encoding framework, the improvement in compression ratio is limited. Neural-enhanced streaming [50, 49, 26, 52, 56] transmits distorted low-bitrate videos and then restores high-fidelity details using post-processing algorithms such as super-resolution. However, these post-processing algorithms rely on prior knowledge learned from training sets. Due to the domain gap, their performance will degrade in unseen scenes [48]. Generative streaming [24, 44, 42, 29] transmits small-sized semantic information, and the receiver can generate video based on these semantics. However, most generative algorithms are designed for specific tasks (e.g., talking video generation driven by facial keypoints [24, 44]), limiting their generality.
@@
With the popularity of AIGC (AI-generated content), Stable Diffusion [38, 10, 9] has attracted extensive attention thanks to its powerful text-to-image generation capability. By pre-training on an internet-scale dataset [6], Stable Diffusion learns prior knowledge of all human visual domains, and simultaneously learns the mapping from text to images. Therefore, Stable Diffusion can generate high-fidelity images based on a brief prompt composed of a few words.
@@
We are motivated by this question: is it possible for Stable Diffusion to replace the video codecs? During streaming, the sender streams prompts instead of streaming encoded videos and the receiver generates videos instead of decoding videos. In this way, the traffic of network video is reduced from the video scale to the text scale, achieving communication efficiency beyond the Shannon limit.
In this paper, we propose Promptus, a system that inverts video frames into Stable Diffusion prompts, achieving ultra-low bitrate video streaming. To bring this vision to fruition, we address the following technical challenges:
@@
First, how to ensure pixel alignment between the generated frames and the real frames. To invert a frame into a prompt, the most straightforward and powerful approach is to manually write a textual description. However, the frame generated using this description can only guarantee semantic consistency with the original frame, while the differences at the pixel level are often substantial, as shown in Figure 1. To achieve pixel alignment, Promptus proposes a gradient descent-based prompt fitting framework. Specifically, the prompt is randomly initialized and then used to generate frame. The pixel-wise loss between the generated frame and the real frame is calculated. The partial derivative of the loss with respect to the prompt will be calculated and the prompt is iteratively optimized using gradient descent. To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid higher-order derivatives. Second, Promptus uses embeddings as prompts instead of text to avoid non-differentiability. Third, Promptus uses a noisy previous frame instead of random noise to reduce latent space distance. Fourth, Promptus combines reconstruction and perceptual loss to enhance the perceptual quality of the generated frames.
@@
Second, how to control the bitrate of the prompt. Since Promptus uses embeddings as prompts, which are matrices with fixed dimensions, the bitrate of the prompt cannot adapt to dynamic network bandwidth. To address this, Promptus proposes a low-rank bitrate control algorithm. Specifically, Promptus integrates the inverse process of low-rank decomposition into the gradient descent, directly fitting the decomposed prompt. The rank is used to control the trade-off between the quality and bitrate of the prompt. When the rank is higher, the representational capability of the prompt is better, and it can describe more details in the frame, but the data size is also larger. Therefore, Promptus adaptively selects the prompt rank based on the currently available bandwidth.
@@
Third, how to perform inter-frame compression on prompts. Promptus inverts each frame into an independent prompt without considering the correlation between prompts across frames. Therefore, when streaming video, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate. To address this, Promptus adds a temporal smoothing regularization during prompt fitting, ensuring that temporally close frames are also sufficiently close in the prompt space. With this, Promptus only needs to sparsely transmit prompts for a few keyframes, while the prompts for the remaining frames can be approximated through linear interpolation of the keyframe prompts.
@@
We evaluated Promptus on test videos from different domains and under real network traces. The results show that: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE [27, 38] and H.265 [5], respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
@@
The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming (§3). (2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion for the first time (§3.1). (3) We build a video streaming system based on Promptus (§4).
2 Motivation and related work
@@ 2.1 Video codec and streaming
Traditional codecs. In network video traffic, most of the content is encoded using traditional codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5]. These traditional video codecs achieve compression primarily by eliminating redundancy in the video. Specifically, the codecs exploit the correlation between adjacent pixel blocks to reduce spatial redundancy in intra-frame prediction. Besides, considering the similarity between consecutive video frames, the codecs employ inter-frame prediction techniques to eliminate temporal redundancy. For moving objects, the codecs perform motion estimation and compensation to further reduce redundancy. Although these techniques achieve efficient compression, the compression ratio has an upper limit due to the finite amount of redundancy. Further compression inevitably requires discarding some non-redundant information, resulting in a drastic decrease in video quality. For example, as shown in Figure 11, when the bitrate is extremely low, videos compressed using H.265 exhibit significant blurriness.
@@
Neural-embedded codecs. With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks [31, 30, 19, 37, 47]. Compared to handcrafted modules, these embedded neural networks can learn more complex and nonlinear motion patterns and intra-frame correlations, thereby achieving lower distortion at the same bitrate. Furthermore, due to the ability to optimize end-to-end, neural-embedded codecs can be trained for tasks beyond compression, such as loss resilience. However, since neural-embedded codecs still adhere to the traditional coding framework, aiming to fully remove intra-frame or inter-frame redundancy, the improvement in compression ratio is limited.
@@
Neural-enhanced streaming. In contrast to reducing redundancy, neural-enhanced streaming [50, 49, 26, 52, 56] actively discards most of the information (including non-redundant information) and transmits highly distorted low-bitrate videos. Then, at the receiver side, neural network-based post-processing algorithms (such as super-resolution) are used to restore high-fidelity details. Neural-enhanced streaming leverages the powerful image restoration and enhancement capabilities of neural networks, effectively recovering lost textures and details, thus ensuring video quality while achieving substantial compression. However, these post-processing algorithms rely on prior knowledge learned from training datasets, such as texture features and edge structures. Due to the complex and diverse nature of real-world video, there exists domain gaps with the training set. The performance of these algorithms often degrades in unseen scenarios [48].
@@
Generative streaming. Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios [24, 44, 42, 29]. For example, in the context of video conferencing, Face-vid2vid [44, 24] extracts facial keypoints in real-time at the sender side for transmission. At the receiver side, it generates dynamic facial videos using the keypoints and static facial images. The bitrate of facial keypoints is much lower than that of videos, thus compressing video conferences to extremely low bitrates. Similarly, Gemino [42] transmits video streams at extremely low resolution. At the receiver side, it estimates facial motion fields from the low-resolution video, then utilizes the motion fields and high-resolution facial images to generate high-resolution facial videos. Instead of driving static images, Reparo [29] proposes a token-based generative streaming. It first uses VQGAN [20] to train a codebook of facial visual features, then maps the video to latent variables using VAE [27], and finally quantizes the latent variables into tokens according to the codebook. Since tokens are indices of the codebook, the bitrate is extremely low. In general, generative streaming can greatly compress videos, but it is often designed for specific tasks (such as video conferencing) and lacks generality.
@@ 2.2 Stable diffusion
Stable Diffusion [38, 10, 9] is the most popular open-source text-to-image generative model. It learns a denoising process from 5.85 billion image-text pairs [6], enabling it to generate high-quality images by denoising the pure noise images. Different noise images lead to different generated images. Since the noise images are randomly sampled from Gaussian distribution, the images generated by Stable Diffusion are random and uncontrollable. Therefore, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts. Specifically, as shown in Figure 2, $p$ represents the user-input prompt. Stable Diffusion first performs word embedding and semantic extraction using the CLIP model [36], converting discrete natural language into a continuous text embedding $c$ (an $m*n$ matrix). The text embeddings and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$. Next, $N^{\prime}$ is input into the denoising process again for $T$ iterations to obtain the denoised $Z$. Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder [27] to generate the image $x$ in the pixel space.
@@
  • Stable Diffusion can effectively reduce the bandwidth overhead of videos. It can generate high-quality images using brief prompts. As shown in Figure 2, a prompt consisting of only a few words can enable Stable Diffusion to generate images with high resolution and complex textures. In codec-based streaming, such high-quality images require high bitrates. Instead of transmitting image, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
@@
  • Stable Diffusion is generalizable to most video domains. Trained on an internet-scale dataset [6] (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet. Consequently, it does not suffer from the domain gap issue and can generate images of any style and content. This gives Stable Diffusion the potential to become a versatile paradigm for video streaming.
@@
  • Stable Diffusion can generate images in real-time. With ongoing advancements in Stable Diffusion, variants capable of real-time image generation have emerged. For instance, StreamDiffusion [28] can generate images at a speed of 100 FPS. This enables Stable Diffusion to meet the real-time requirements of video streaming.
@@
However, due to the inability to precisely control image generation, Stable Diffusion cannot meet the fidelity requirements of video streaming. Specifically, Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images. The most straightforward solution is to automatically extract text descriptions for target images using Image/Video Captioning algorithms [46, 23] and use these extracted descriptions as prompts to generate images. Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images lacking many details or having large structural differences with the target images. In fact, even if human intelligence is used to manually describe images, it is impossible to generate images that are pixel-aligned with the target images, as shown in Figure 1. In addition to using text as prompts, ControlNet [54] can utilize images (such as masks) as prompts, controlling Stable Diffusion to generate images that align with the contours of the prompt image. However, apart from the contours, details such as color and texture cannot be aligned. In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original frames at the pixel level, making them unsuitable for video streaming.
@@
To materialize the above potential benefits, Promptus is the first to successfully invert frames into prompts while ensuring pixel-level alignment. Promptus is related to some work on Text Inversion [21, 39, 25], both of which learn prompts from images in an inverse manner. However, Text Inversion aims to learn new words that represent the appearance of specific objects from images, aligning only at the semantic level. Text Inversion has also been used for image compression [35]. But similarly, the inverse prompt is only used for semantic alignment, while pixel alignment is achieved by using low-resolution images as conditions.

### v2

@@ Abstract
With the exponential growth of video traffic, traditional video streaming systems are approaching their limits in compression efficiency and communication capacity. To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive semantic communication system that streaming prompts instead of video content, which represents real-world video frames with a series of "prompts" for delivery and employs Stable Diffusion to generate videos at the receiver. To ensure that the prompt representation is pixel-aligned with the original video, a gradient descent-based prompt fitting framework is proposed. Further, a low-rank decomposition-based bitrate control algorithm is introduced to achieve adaptive bitrate. For inter-frame compression, an interpolation-aware fitting algorithm is proposed. Evaluations across various video genres demonstrate that, compared to H.265, Promptus can achieve more than a 4x bandwidth reduction while preserving the same perceptual quality. On the other hand, at extremely low bitrates, Promptus can enhance the perceptual quality by 0.139 and 0.118 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Our work opens up a new paradigm for efficient video communication. Promptus is open-sourced at: https://github.com/JiangkaiWu/Promptus.
@@ 1 Introduction
With the rapid development of streaming applications (such as Youtube, Netflix), the traffic of network video has been continuously growing. To reduce traffic, video codecs represented by VP8 (bankoski2011technical, ), VP9 (mukherjee2015technical, ), H.264 (264, ) and H.265 (265, ) are widely used to compress videos. These codecs achieve compression by removing spatial and temporal redundancies. However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit (shannon1948mathematical, )). In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts. In recent years, some deep learning-based codecs (lu2019dvc, ; lin2020m, ; djelouah2019neural, ) and streaming frameworks (zhou2022cadm, ; jiang2022wireless, ; sivaraman2024gemino, ; li2023reparo, ) have been proposed to improve compression ratio, but they are limited either by performance or by generality.
@@
Figure 1: Promptus can invert a given image into a prompt. Based on this prompt, Stable Diffusion can generate an almost identical image to the original. In contrast, existing methods can only generate semantically similar images, while the differences at the pixel level are substantial. In this way, Promptus streams prompts instead of videos, significantly reducing bandwidth overhead.
@@
With the popularity of generative AI, Stable Diffusion (rombach2022high, ; sd, ) has attracted extensive attention thanks to its powerful text-to-image generation capability. By pre-training on an internet-scale dataset (LAION, ), Stable Diffusion learns prior knowledge of nearly all human visual domains, and simultaneously learns the mapping from text to images. Therefore, Stable Diffusion can generate high-fidelity images based on a brief prompt composed of a few words.
@@
We are motivated by this question: is it possible for Stable Diffusion to replace the video codecs? During streaming, the sender streams prompts instead of streaming encoded videos and the receiver generates videos instead of decoding videos. In this way, the traffic of network video is reduced from the video scale to the text scale, greatly improving video communication efficiency. In this paper, we propose Promptus, a system that represents real-world video frames with Stable Diffusion prompts, achieving ultra-low bitrate video streaming. To bring this vision to fruition, we address the following technical challenges:
@@
First, how to ensure pixel alignment between the generated frames and the real frames. To invert a frame into a prompt, the most straightforward and powerful approach is to manually write a textual description. However, the frame generated using this description can only guarantee semantic consistency with the original frame, while the differences at the pixel level are often substantial, as shown in Figure 1. To achieve pixel alignment, Promptus proposes a gradient descent-based prompt fitting framework. Specifically, the prompt is randomly initialized and then used to generate frame. The pixel-wise loss between the generated frame and the real frame is calculated. The partial derivative of the loss with respect to the prompt will be calculated and the prompt is iteratively optimized using gradient descent. To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid unrolled long gradient chains. Second, Promptus uses embeddings as prompts instead of text to avoid non-differentiability. Third, Promptus uses a noisy previous frame instead of random noise to reduce latent space distance. Fourth, Promptus combines reconstruction and perceptual loss to enhance the perceptual quality of the generated frames.
@@
Second, how to control the bitrate of the prompt. Since Promptus uses embeddings as prompts, which are matrices with fixed dimensions, the bitrate of the prompt cannot adapt to dynamic network bandwidth. To address this, Promptus proposes a low-rank bitrate control algorithm. Specifically, Promptus integrates the inverse process of low-rank decomposition into the gradient descent, directly fitting the decomposed prompt. The rank is used to control the trade-off between the quality and bitrate of the prompt. When the rank is higher, the representational capability of the prompt is better, and it can describe more details in the frame, but the data size is also larger. So Promptus adaptively selects the prompt rank based on the currently available bandwidth.
@@
Third, how to perform inter-frame compression on prompts. Promptus inverts each frame into an independent prompt without considering the correlation across frames. Therefore, when streaming, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate. To address this, Promptus adds an interpolation-aware fitting algorithm, ensuring that consecutive frames also exhibit continuity in the prompt space. With this, Promptus only needs to sparsely transmit prompts for a few keyframes, while the prompts for the remaining frames can be approximated through linear interpolation of the keyframe prompts.
@@
We evaluated Promptus on test videos from different domains and under real-world network traces. The results show that: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, compared to H.265 (265, ), Promptus provides more than a 4x bandwidth reduction while preserving the same perceptual quality. On the other hand, at extremely low bitrates, Promptus can enhance the perceptual quality by 0.139 and 0.118 (in LPIPS (zhang2018unreasonable, )) compared to VAE (kingma2014auto, ; rombach2022high, ) and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%.
@@
The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming. (2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion. (3) We build a video streaming system based on prompts, significantly boosting video communication efficiency.
2 Related work and Background
@@ 2.1 Video codec and streaming
Video codecs. In network video traffic, most of the content is encoded using traditional codecs represented by VP8 (bankoski2011technical, ), VP9 (mukherjee2015technical, ), H.264 (264, ) and H.265 (265, ). These traditional video codecs achieve compression primarily by eliminating redundancy in the video. Specifically, the codecs exploit the correlation between adjacent pixel blocks to reduce spatial redundancy in intra-frame prediction. Besides, considering the similarity between consecutive video frames, the codecs employ inter-frame prediction techniques to eliminate temporal redundancy. Although these techniques achieve efficient compression, the compression ratio has an upper limit due to the finite amount of redundancy. With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks (lu2019dvc, ; lin2020m, ; djelouah2019neural, ; rippel2019learned, ; li2024neural, ; sheng2024prediction, ). Compared to handcrafted modules, these embedded neural networks can learn more complex and nonlinear motion patterns and intra-frame correlations, thereby achieving lower distortion at the same bitrate. However, neural-embedded codecs still adhere to the traditional coding framework (aiming to remove redundancy), so the improvement is limited.
@@
Neural-enhanced streaming. In contrast to reducing redundancy, neural-enhanced streaming (park2023omnilive, ; zhou2022cadm, ) actively discards most of the information (including non-redundant information), transmitting highly distorted low-bitrate videos. Then, at the receiver, neural network-based post-processing methods (such as super-resolution) are used to restore high-fidelity details. Thanks to the powerful image restoration capabilities of neural networks, the lost details can be effectively recovered, thus ensuring video quality while achieving substantial compression. However, these post-processing neural networks rely on prior knowledge learned from training datasets. Due to the complex and diverse nature of real-world video, there exists domain gaps with the training set. The performance often degrades in unseen scenarios (yang2021real, ). A feasible solution is to fine-tune the post-processing neural networks for each video segment and send the fine-tuned neural networks along with the video, like NAS (yeo2017will, ; yeo2018neural, ; yeo2022neuroscaler, ; yeo2020nemo, ; kim2020neural, ; zhang2022yuzu, ). However, the transmission of the neural networks inevitably increases the overall bitrate. Especially for shorter videos (like Tiktok), the overhead of neural networks is unacceptable.
@@
Generative streaming. Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios (jiang2022wireless, ; wang2021one, ; sivaraman2024gemino, ; li2023reparo, ). For example, in the context of video conferencing, Face-vid2vid (wang2021one, ; jiang2022wireless, ) extracts facial keypoints in real-time at the sender side for transmission. At the receiver side, it generates dynamic facial videos using the keypoints and static facial images. The bitrate of facial keypoints is much lower than that of videos, thus compressing video conferences to extremely low bitrates. Similarly, Gemino (sivaraman2024gemino, ) transmits video streams at extremely low resolution. At the receiver side, it estimates facial motion fields from the low-resolution video, then utilizes the motion fields and high-resolution facial images to generate high-resolution facial videos. Instead of driving static images, Reparo (li2023reparo, ) proposes a token-based generative streaming. It first uses VQGAN (esser2021taming, ) to train a codebook of facial visual features, then maps the video to latent variables using VAE (kingma2014auto, ), and finally quantizes the latent variables into token indices according to the codebook. Since only indices need to be transmitted, the bitrate is very low. In general, generative streaming can greatly compress videos, but it is often designed for specific tasks (such as video conferencing) and lacks generality.
@@ 2.2 Stable diffusion
Stable Diffusion (rombach2022high, ; sd, ) is the most popular open-source text-to-image generative model. It learns a denoising process from 5.85 billion image-text pairs (LAION, ), enabling it to generate high-quality images by denoising the pure noise images. Different noise images lead to different generated images. Since the noise images are randomly sampled from Gaussian distribution, the generated images are random and uncontrollable. Thus, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts. Specifically, as shown in Figure 2, let $p$ represent the user-input prompt. Stable Diffusion first performs word embedding and semantic extraction using the CLIP model (radford2021learning, ), converting discrete natural language $p$ into a continuous text embedding $c$ (an $m*n$ matrix). The text embeddings $c$ and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$. Next, $N^{\prime}$ is input into the denoising process again for $T$ iterations to obtain the denoised $Z$. Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder (kingma2014auto, ) to generate the image $x$ in the pixel space.
@@
  • Stable Diffusion can effectively reduce the bandwidth overhead of videos. It can generate high-quality images using brief prompts. As shown in Figure 2, a prompt consisting of only a few words can enable Stable Diffusion to generate images with high resolution and complex textures. In codec-based streaming, such high-quality images require high bitrates. Instead of transmitting images, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
  • Stable Diffusion decouples the prompt size from the video resolution. A key insight in Stable Diffusion’s image generation is that the resolution of the generated images is independent of the prompt size. As illustrated in Figure 2, since the denoising process does not change the resolution, the resolution of the generated image $x$ is solely dependent on the resolution of the input noise image $N$. Since the noise image $N$ is simply obtained through random sampling from a Gaussian distribution, the resolution can be set freely. Given the same prompt, using noise images $N$ sampled at different resolutions will generate high-quality images that are semantically similar but differ in resolution. In contrast, encoder-based compression (such as VAE (kingma2014auto, ) and H.265) inevitably requires an increase in data size as the resolution rises. This is the fundamental difference between prompts and encoders and is the key to achieving ultra-low bitrate video communication.
@@
  • Stable Diffusion is generalizable to most video domains. Trained on an internet-scale dataset LAION (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet. Consequently, it does not suffer from the domain gap issue and can generate images of any style and content. This gives Stable Diffusion the potential to become a versatile paradigm for video streaming.
@@
  • Stable Diffusion can generate images in real-time. With ongoing advancements in Stable Diffusion, variants capable of real-time image generation have emerged. For instance, StreamDiffusion kodaira2023streamdiffusion can generate images at a speed of 100 FPS. This enables Stable Diffusion to meet the real-time playback requirements at the receiver.
@@
However, due to the inability to precisely control image generation, Stable Diffusion cannot meet the fidelity requirements of video streaming. Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images. The most straightforward solution is to extract text descriptions for target images using Image/Video Captioning algorithms (yang2023vid2seq, ; hu2022scaling, ) and use these extracted descriptions as prompts to generate images. Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images having large structural differences with the target images. In fact, even if human intelligence is used to manually describe images, it is impossible to generate images that are pixel-aligned with the target images, as shown in Figure 1. Besides text, ControlNet (zhang2023adding, ) can also use images (such as masks and edges) as prompts, controlling Stable Diffusion to generate images that align with the contours of the image prompts. However, apart from the contours, details such as color and texture cannot be aligned qiao2024latency . In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original ones at the pixel level, making them unsuitable for video streaming.
@@
To materialize the above potential benefits, Promptus is an initial effort to invert frames into prompts while ensuring pixel-level alignment. It is related to some works on Text Inversion (gal2022image, ; ruiz2023dreambooth, ; kawar2023imagic, ), both of which learn prompts from images in an inverse manner. However, Text Inversion aims to learn text embeddings that represent the appearance of specific objects from multiple images, aligning only at the semantic level. Text Inversion has also been used for image compression (pan2022extreme, ). But similarly, the inverse prompt is only used for semantic alignment, while pixel alignment is achieved by using low-resolution images as conditions.